Let a 4x4 matrix V =

1 x x^2 x^3 //

1 y y^2 y^3 //

1 z z^2 z^3 //

1 w w^2 w^3


Prove that det(V)=(y-x)(z-x)(w-x)(z-y)(w-y)(w-z)


1
Expert's answer
2020-02-17T08:30:56-0500

"Det(V)=\begin{vmatrix} 1 & x & x^2 & x^3\\ 1 & y & y^2 & y^3\\ 1 & z & z^2 & z^3\\ 1 & w & w^2 & w^3\\ \end{vmatrix}"

"=\begin{vmatrix} 1 & x & x^2 & x^3\\ 0 & y-x & y^2-x^2 & y^3-x^3\\ 0 & z-x & z^2-x^2 & z^3-x^3\\ 0 & w-x & w^2-x^2 & w^3-x^3\\ \end{vmatrix}"

Row 2 = Row 2 - Row 1; Row 3 = Row 3 - Row 1; Row 4 = Row 4 - Row 1

"=\begin{vmatrix} 
1 & x & x^2 & x^3\\ 
0 & y-x & (y-x)(y+x) & (y-x)(y^2+yx+x^2)\\ 
0 & z-x & (z-x)(z+x) & (z-x)(z^2+zx+x^2)\\ 
0 & w-x & (w-x)(w+x) & (w-x)(w^2+wx+x^2)\\ 
\end{vmatrix}"

"=(y-x)(z-x)(w-x) 
\begin{vmatrix} 
1 & x & x^2 & x^3\\ 
0 & 1 & (y+x) & (y^2+yx+x^2)\\ 
0 & 1 & (z+x) & (z^2+zx+x^2)\\ 
0 & 1 & (w+x) & (w^2+wx+x^2)\\ 
\end{vmatrix}"

By taking out "(y-x)" from Row 2, "(z-x)" from Row 3, "(w-x)" from Row 4

"=(y-x)(z-x)(w-x) 
\begin{vmatrix} 
1 & x & x^2 & x^3\\ 
0 & 1 & y+x & (y^2+yx+x^2)\\ 
0 & 0 & z-y & (z^2-y^2+zx-yx)\\ 
0 & 0 & w-y & (w^2-y^2+wx-yx)\\ 
\end{vmatrix}"

Row 3 = Row 3 - Row 2; Row 4 = Row 4 - Row 2

"=(y-x)(z-x)(w-x) \n\\begin{vmatrix} \n1 & x & x^2 & x^3\\\\ \n0 & 1 & y+x & (y^2+yx+x^2)\\\\ \n0 & 0 & z-y & ((z-y)(z+y)+(z-y)x)\\\\ \n0 & 0 & w-y & ((w-y)(w+y)+(w-y)x)\\\\ \n\\end{vmatrix}\u200b\\\\\n=(y-x)(z-x)(w-x) \n\\begin{vmatrix} \n1 & x & x^2 & x^3\\\\ \n0 & 1 & y+x & (y^2+yx+x^2)\\\\ \n0 & 0 & z-y & (z-y)(z+y+x)\\\\ \n0 & 0 & w-y & (w-y)(w+y+x)\\\\ \n\\end{vmatrix}"

"=(y-x)(z-x)(w-x)(z-y)(w-y) 
\begin{vmatrix} 
1 & x & x^2 & x^3\\ 
0 & 1 & y+x & (y^2+yx+x^2)\\ 
0 & 0 & 1 & (z+y+x)\\ 
0 & 0 & 1 & (w+y+x)\\ 
\end{vmatrix}"

By taking out "(z-y)" from Row 3, "(w-y)" from Row 4

"=(y-x)(z-x)(w-x)(z-y)(w-y) 
\begin{vmatrix} 
1 & x & x^2 & x^3\\ 
0 & 1 & y+x & (y^2+yx+x^2)\\ 
0 & 0 & 1 & (z+y+x)\\ 
0 & 0 & 0 & (w-z)\\ 
\end{vmatrix}"

Row 4 = Row 4 - Row 3

"=(y-x)(z-x)(w-x)(z-y)(w-y)(w-z) 
\begin{vmatrix} 
1 & x & x^2 & x^3\\ 
0 & 1 & y+x & (y^2+yx+x^2)\\ 
0 & 0 & 1 & (z+y+x)\\ 
0 & 0 & 0 & 1\\ 
\end{vmatrix}"

By taking out "(w-z)" from Row 4

"=(y-x)(z-x)(w-x)(z-y)(w-y)(w-z)."

Since "\begin{vmatrix} 1 & x & x^2 & x^3\\ 0 & 1 & y+x & (y^2+yx+x^2)\\ 0 & 0 & 1 & (z+y+x)\\ 0 & 0 & 0 & 1\\ \end{vmatrix}
=1"

Hence proved.